1.) How many different ways can 3 cards be drawn from a deck of 52
cards without replacement?
2.) How many different ways can 3 cards be drawn from a deck of 52
cards with replacement?
3.) A corporation has 5 officers to choose from which 3 are selected
to comprise the board of directors. How many combinations are there?
(Order is not important)
4.) Convert to a percent in simplest form: 0.06

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

This is 52*51*50=132,600
This would be 52*52*52=140,600
This is 5C3=5!/3!2!=10
6%. Move the decimal two places to the right.

Answer by ikleyn(52781)

.
The correct answer to the second question is THIS

    52*52*52 = 140608.